If water weighs 1000 kg per cubic meter, determine the total weight of the water that a 1 -meter diameter cylindrical tank carries. *





1
Expert's answer
2021-04-29T10:42:28-0400

The volume of 1-meter diameter cylindrical tank 1 m tall is


"V=Ah=\pi r^2h=3.14·0.5^2·1=0.785\text{ m}^3."

The mass is


"m=\rho V=1000·0.785=785\text{ kg}."